Can you please help me solve this I keep getting the operations mixed up:
Method of Substitution
5a = 3 - b
5a = 4 - b
Thank you

Answer by Alan3354(69443)

Method of Substitution
5a = 3 - b
5a = 4 - b
-----------
Since the both equal 5a:
3 - b = 4- b
Add b
3 = 4
----------
The system is inconsistent.
No solution.